# The local moduli of Sasaki-Einstein $\mathbb{Q}$-homology 7-spheres and invertible polynomials

### v1

@@ 3.2. $\mathbb{Q}$-homology 7-spheres: Thom-Sebastiani sums of invertible polynomials
Remark 3.2.5.
We have some comments.
  • a)
    When $f$ is a polynomial of type II such that its associated weight vector $\bf w$ admits a polynomial of type I, then part a) of the previous theorem holds. Thus, $\mu=m_{2}-3$.
  • b)
    If $f$ is a polynomial of type III whose associated weight vector $\bf w$ admits a polynomial of type II but not type I, then we can use part b) of the previous theorem. Thus, $\mu=\dfrac{m_{2}}{v_{1}}-\dfrac{1}{v_{1}}-2.$
  • c)
    Finally, if $f$ is a polynomial of type III whose associated weight vector admits polynomial of type I and II, then $\mu=\dfrac{m_{2}}{v_{0}v_{1}}-\dfrac{1}{v_{0}}-\dfrac{1}{v_{1}}-1.$
3.3. $\mathbb{Q}$-homology sphere: the Berglund-Hübsch dual of chain-cycle polynomials
As mentioned before, in [14] it was proven that the Berglund-Hübsch transpose rule only produces twins for singularities of cycle type or of type I and type III. Actually the types are preserved under the Berglund-Hübsch transpose:
@@ 3.3. $\mathbb{Q}$-homology sphere: the Berglund-Hübsch dual of chain-cycle polynomials
From Theorem 3.2.1 the real dimension $\mu_{\mathbb{R}}$ of the local moduli for links arising from polynomials of type I and type III are given by $\mu_{\mathbb{R}}=2[\frac{m_{2}}{v_{0}v_{1}}-\frac{1}{v_{0}}-\frac{1}{v_{1}}-1]$, and from Theorem 3.1.1 $\mu_{\mathbb{R}}=0$ in case the singularity is given by a cycle type polynomial given in Subsection 3.1. Thus we only need to study polynomials of type II, that is polynomials of chain-cycle type, where the Berglund-Hübsch transpose rule does not preserve neither torsion nor Milnor number, and hence $m_{2}$, $v_{0}$ and $v_{1}$ vary.
@@
The complex dimension of the local moduli for $Z_{f^{T}}$ is given by
$\dim H^{0}(\mathbb{P}(\tilde{\bf w}),\mathcal{O}(\tilde{d}))-\sum_{i}\dim H^{0}(\mathbb{P}(\tilde{\bf w}),\mathcal{O}(\tilde{w}_{i}))$ (3.3.6)
We begin computing $\dim H^{0}(\mathbb{P}(\tilde{\bf w}),\mathcal{O}(\tilde{d}))$. Since this refers to the number of all possible monomials $z_{0}^{x_{0}}z_{1}^{x_{1}}z_{2}^{x_{2}}z_{3}^{x_{3}}z_{4}^{x_{4}}$ of degree $\tilde{d}$, we calculate this number counting the solutions of the Diophantine equation
@@
where the unknowns $x_{i}$ are non-negative integers such that at least one of them is non-zero. We will solve this equation adding certain mild constraints on the exponent $a_{1}$ and $m_{3}$. As we will see the remaining cases of interest, those whose corresponding links admit Sasaki-Einstein metrics, can be computed case by case.
@@
Let $f^{T}$ be an invertible polynomial described as in (3.3.3) with an associated weight vector $\tilde{\bf w}$ given as in (3.3.5). If we add the additional conditions $\gcd(a_{1},m_{3})=1$ and $a_{1}>2$, then the equation (3.3.7) has exactly five solutions. Moreover, the set of generators of $H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(\tilde{d}))$ is given by the set
@@
    Since $\gcd(a_{1},m_{3})=1$, then we have $m_{3}\mid(m_{3}-\tilde{v}_{2}x_{2}-\tilde{v}_{3}x_{3}-\tilde{v}_{4}x_{4})$. That is, the expression $\tilde{v}_{2}x_{2}+\tilde{v}_{3}x_{3}+\tilde{v}_{4}x_{4}$ can assume two values: $m_{3}$ or 00. First, we suppose that $\tilde{v}_{2}x_{2}+\tilde{v}_{3}x_{3}+\tilde{v}_{4}x_{4}=m_{3}$. This forces $x_{1}=0$. Moreover, since $\tilde{v}_{2}$, $\tilde{v}_{3}$ and $\tilde{v}_{4}$ are given in (3.3.4) and these describe the weight vector $\tilde{\bf w}$ in (3.3.5), then the equation $\tilde{v}_{2}x_{2}+\tilde{v}_{3}x_{3}+\tilde{v}_{4}x_{4}=m_{3}$ has exactly three solutions: $(a_{2},1,0),(0,a_{3},1)$ and $(1,0,a_{4})$. Thus, we obtain three solutions for the equation (3.3.8): $(0,0,a_{2},1,0),(0,0,0,a_{3},1)$ and $(0,0,1,0,a_{4})$. Finally, if we suppose that $\tilde{v}_{2}x_{2}+\tilde{v}_{3}x_{3}+\tilde{v}_{4}x_{4}=0$, then $x_{1}=a_{1}$. In this case the only solution is $(0,a_{1},0,0,0)$.
@@
Let $f$ be and invertible polynomial as in (3.3.1) with weights satisfying conditions (3.2.1). If its Berglund-Hübsch transpose $f^{T}$ with associated weight vector $\bf\tilde{w}$ satisfies the conditions given in the Lemma 3.3.1. Then the complex dimension of the local moduli of the orbifold $Z_{f^{T}}=\{f^{T}=0\}/\mathbb{C}^{*}(\bf{\tilde{w}})$ is zero. Moreover, each $H^{0}(\mathbb{P}(\tilde{\bf w}),\mathcal{O}(\tilde{w}_{i}))$ is generated by $z_{i}$ for $i=0,\ldots,4.$
@@
This result follows from Lemma 3.3.1 and the fact that $\dim H^{0}(\mathbb{P}(\tilde{\bf w}),\mathcal{O}(\tilde{w}_{i}))\geqslant 1$ for each $i$ and $\sum_{i=1}^{4}H^{0}(\mathbb{P}(\tilde{\bf w}),\mathcal{O}(\tilde{w}_{i}))\leqslant 5.$
@@
Let $f^{T}$ be an invertible polynomial described in (3.3.3) with associated weight vector $\tilde{\bf w}$ as in (3.3.5). If we add the additional condition $a_{1}=2$, then Equation (3.3.7) has six solutions. Moreover, the set of generators of $H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(\tilde{d}))$ is given by the set
@@
As $\gcd(m_{2},m_{3})=1$, then $m_{2}\mid(x_{0}+m_{2}x_{1})$. This implies that $m_{2}\mid x_{0}$. Moreover, from (3.3.9) we have $m_{3}x_{0}\leqslant 2m_{3}m_{2}$. Since $m_{3}\neq 0$, then $x_{0}\leqslant 2m_{2}$. Thus, $x_{0}$ can take three values: 00, $m_{2}$ or $2m_{2}$.
@@
        Nevertheless, the solutions of this last equation are $(a_{2},1,0),(0,a_{3},1)$ and $(1,0,a_{4})$, which implies that $12\hat{x}_{i}=$ 1 for any $i$. That is, $\hat{x}_{i}$ is not integer, which is not possible.
@@
        which has trivial solution. Therefore, the solutions of Equation (3.3.9) is $(0,2,0,0,0)$.
@@
Let $f$ be and invertible polynomial as in (3.3.1) with weights satisfying conditions (3.2.1). If its Berglund-Hübsch transpose $f^{T}$ with associated weight vector $\bf\tilde{w}$ satisfies the conditions given in the Lemma 3.3.2. Then the complex dimension of the local moduli of the orbifold $Z_{f^{T}}=\{f^{T}=0\}/\mathbb{C}^{*}(\bf{\tilde{w}})$ is zero. Moreover, $H^{0}(\mathbb{P}(\tilde{\bf w}),\mathcal{O}(\tilde{w}_{1}))$ is generated by $z_{0}^{m_{2}}$ and $z_{1}$ while for $i\not=1$ $H^{0}(\mathbb{P}(\tilde{\bf w}),\mathcal{O}(\tilde{w}_{i}))$ is generated by $z_{i}.$
@@
First, we will prove that $\dim H^{0}(\mathbb{P}(\tilde{\bf w}),\mathcal{O}(\tilde{w}_{1}))\geqslant 2$. In fact, it is equivalent to show that the quantity of solutions of the following Diophantine equation is not less that two:
@@
Here, we can exhibit at least two solutions: $(m_{2},0,0,0,0)$ and $(0,1,0,0,0)$. Therefore, we obtain $\dim H^{0}(\mathbb{P}(\tilde{\bf w}),\mathcal{O}(\tilde{w}_{1}))\geqslant 2$.
On the other hand, we know that $\dim H^{0}(\mathbb{P}(\tilde{\bf w}),\mathcal{O}(\tilde{w}_{i}))\geqslant 1$ for $i\neq 2$. This implies that
$66\leqslant\sum_{i=0}^{4}\dim H^{0}(\mathbb{P}(\tilde{\bf w}),\mathcal{O}(\tilde{w}_{i}))\leqslant\dim H^{0}(\mathbb{P}(\tilde{\bf w}),\mathcal{O}(\tilde{d}))=$ 6 .
@@
$\dim H^{0}(\mathbb{P}(\tilde{\bf w}),\mathcal{O}(\tilde{d}))-\sum_{i=0}^{4}\dim H^{0}(\mathbb{P}(\tilde{\bf w}),\mathcal{O}(\tilde{w}_{i}))=0.$
@@
Remark 3.3.1.
For $\mathbb{P}(\mathbf{\tilde{w}})=\operatorname{Proj}(S(\mathbf{\tilde{w}}))$, the complex automorphisms of the graded ring $S(\mathbf{\tilde{w}})$ with can be defined on generators for the two types of polynomials given in this subsection thanks to Propositions 3.3.1 and 3.3.2. Notice that the generating set of $H^{0}(\mathbb{P}(\mathbf{\tilde{w}}),\mathcal{O}({\tilde{d}}))$ in all these cases is given by quasi-smooth monomials as in the previous cases, so we can describe the local moduli of the corresponding orbifold, a $\mathbb{Q}$-homology projective 3-space, as done previously. Indeed, let $\alpha_{i},\beta_{1}\in\mathbb{C}^{*},$ then we have
  • If $f$ is a polynomial as in Lemma 3.3.1, then $G(\mathbf{\tilde{w}})$ is given on generators by
@@
    and the moduli of the orbifold $Z_{f^{T}}$ can be expressed by the quotient
    $Span\{z_{0}^{m_{2}}z_{1},z_{1}^{a_{1}},z_{3}z_{2}^{a_{2}},{z_{4}}z_{3}^{a_{3}},z_{2}z_{4}^{a_{4}}\}/G(\mathbf{\tilde{w}})$
@@
  • If $f$ is a polynomial as in Lemma 3.3.2, then $G(\mathbf{\tilde{w}}))$ is given on generators by
@@
    and the moduli of the orbifold $Z_{f^{T}}$ can be expressed by the quotient
    $Span\left\{z_{0}^{m_{2}}z_{1},z_{0}^{2m_{2}},z_{1}^{2},z_{3}z_{2}^{a_{2}},z_{4}z_{3}^{a_{3}},z_{2}z_{4}^{a_{4}}\right\}/G(\mathbf{\tilde{w}})$
@@
In [14], Theorem 4.1 we established the existence of 75 new $\mathbb{Q}$-homology 7-spheres admitting Sasaki-Einstein metrics with not well-formed quotients, that is, with some group elements having codimension 1 fixed point sets. All these links, which are listed in a table in the Appendix in [14], are given by polynomials of chain-cycle type of type II with weights satisfying the conditions given in (3.2.1). So we have:
Theorem 3.3.1.
Let $f$ be and invertible polynomial as in (3.3.1) with weights satisfying conditions (3.2.1). If its Berglund-Hübsch transpose $f^{T}$ with associated weight vector $\bf\tilde{w}$ satisfies the conditions given in either Lemma 3.3.1 or Lemma 3.3.2 and additionally $f$ belongs to one of the 236 $\mathbb{Q}$-homology 7-spheres admitting Sasaki-Einstein metrics found in [5] and [14], then the dimension of the local moduli space of Sasaki-Einstein metrics of the $\mathbb{Q}$-homology 7-spheres at $L_{f^{T}}$ equals zero. Thus $\mathbb{Q}$-homology 7-spheres given as links coming from polynomials $f^{T}$ as above do not admit inequivalent families of Sasaki-Einstein structures.
@@
From the list of 75 new examples of Sasaki-Einstein $\mathbb{Q}$-homology 7-sphere given in the Appendix of [14], there are 7 weight vectors $\bf\tilde{w}$ that do not satisfy the additional conditions of either Lemma 3.3.1 or Lemma 3.3.2, actually all of the elements of this table satisfy $a_{1}>2$ but $\gcd(m_{3},a_{1})\not=1.$ For these, we have the following table where $\mu_{\mathbb{R}}$ denotes the real dimension of the local moduli of Sasaki-Einstein metrics of $\mathbb{Q}$-homology 7-spheres at $L_{f^{T}}.$
$\bf\tilde{w}$ $\tilde{d}$ $m_{3}$ $\dim H^{0}(\mathbb{P}(\tilde{\bf w}),\mathcal{O}(\tilde{d}))$ $\sum_{i=0}^{4}\dim H^{0}(\mathbb{P}(\tilde{\bf w}),\mathcal{O}(\tilde{w}_{i}))$ $\mu_{\mathbb{R}}$

### v2

@@ 3.2. Rational homology 7-spheres: Thom-Sebastiani sums of invertible polynomials
Example 3.2.1.
Let us consider the following examples, where all the polynomials are taken from the list of Johnson and Kollár of anticanonically embedded Fano Kähler-Einstein 3-folds.
  1. (1)
    The polynomial $f=z_{0}^{9}+z_{1}^{9}+z_{4}z_{2}^{2}+z_{2}z_{3}^{2}+z_{3}z_{4}^{19}$ of type I has associated weight vector $\textbf{w}=(77,77,333,180,27)$ and degree $d=m_{3}m_{2}=693$, where $m_{3}=77$ and $m_{2}=9$. The generating set of $H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(d))$ is given by the set of thirteen monomials
    $\{z_{0}^{9},z_{0}^{8}z_{1},\dots,z_{0}z_{1}^{8},z_{1}^{9},z_{4}z_{2}^{2},z_{2}z_{3}^{2},z_{3}z_{4}^{19}\}.$
    Then the moduli of the orbifold $Z_{f}$ is included in
    $Span\left\{z_{0}^{9},z_{0}^{8}z_{1},\dots,z_{0}z_{1}^{8},z_{1}^{9},z_{4}z_{2}^{2},z_{2}z_{3}^{2},z_{3}z_{4}^{19}\right\}/G(\mathbf{w}).$
    where $G(\textbf{w})$ is expressed as above. Thus the complex dimension of this moduli is six and the corresponding link $L_{f}$ has a local moduli of Sasaki-Einstein metrics of real dimension twelve.
  2. (2)
    The polynomial $f=z_{0}^{125}+z_{0}z_{1}^{4}+z_{4}z_{2}^{2}+z_{2}z_{3}^{7}+z_{3}z_{4}^{3}$ of type II with degree $d=m_{3}m_{2}=5375$, where $m_{3}=43$ and $m_{2}=125$, and with associated weight vector is $\textbf{w}=(43,1333,1875,500,1625)$. We notice that $\bf w$ does not admit polynomials of type I. Since $v_{1}=31$, then we have that the generating set of $H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(d))$ is given by $\left(\left\lfloor\frac{125}{31}\right\rfloor+1\right)+3=8$ monomials. Indeed, the generating set is given by
    $\{z_{0}^{125},z_{0}^{94}z_{1},z_{0}^{63}z_{1}^{2},z_{0}^{32}z_{1}^{3},z_{0}z_{1}^{4},z_{4}z_{2}^{2},z_{2}z_{3}^{7},z_{3}z_{4}^{3}\}$
    The moduli of the orbifold $Z_{f}$ is obtained in a similar way as above. Thus the complex dimension of this moduli is two and the corresponding link $L_{f}$ has a local moduli of Sasaki-Einstein metrics of real dimension four.
  3. (3)
    The polynomial $f=z_{1}z_{0}^{5}+z_{0}z_{1}^{15}+z_{4}z_{2}^{2}+z_{2}z_{3}^{4}+z_{3}z_{4}^{4}$ of type III, with associated weight vector $\textbf{w}=(231,66,481,185,259)$ and degree $d=m_{3}m_{2}=1221$, where $m_{3}=33$ and $m_{2}=37$. Here, the weight vector $\bf w$ does not admit polynomials type I or II. Moreover, as $v_{0}=7$ and $v_{1}=2$, then the generating set of $H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(d))$ is formed by $\left(\left\lfloor\frac{37}{(7)(2)}\right\rfloor+1\right)+3=6$ given by
    $\{z_{0}^{5}z_{1},z_{0}^{3}z_{1}^{8},z_{0}z_{1}^{15},z_{4}z_{2}^{2},z_{2}z_{3}^{4},z_{3}z_{4}^{4}\}.$
    It follows that the the complex dimension of this moduli is one and the corresponding link $L_{f}$ has a local moduli of Sasaki-Einstein metrics of real dimension two.
3.3. Rational homology sphere: the Berglund-Hübsch dual of chain-cycle polynomials
As mentioned before, in [16] it was proven that the Berglund-Hübsch transpose rule only produces twins for singularities of cycle type, type I and type III. Actually these types are preserved under the Berglund-Hübsch transpose:
@@ 3.3. Rational homology sphere: the Berglund-Hübsch dual of chain-cycle polynomials
So the the moduli of orbifolds $Z_{f^{T}}$ determined by Berglund-Hübsch transpose dual $f^{T}$ of $f$ for cycle polynomials and polynomials of type I and type III have moduli described by Propositions 3.2.2 and 3.1.1. From Proposition 3.2.2 the real dimension $\mu_{\mathbb{R}}$ of the local moduli for links arising from polynomials of type I and type III are given by $\mu_{\mathbb{R}}=2[\frac{m_{2}}{v_{0}v_{1}}-\frac{1}{v_{0}}-\frac{1}{v_{1}}-1]$, and from Proposition 3.1.2 $\mu_{\mathbb{R}}=0$ in case the singularity is given by a cycle type polynomial given in Subsection 3.1. Thus we only need to study polynomials of type II, that is polynomials of chain-cycle type, where the Berglund-Hübsch transpose rule does not preserve neither torsion nor Milnor number, and hence $m_{2}$, $v_{0}$ and $v_{1}$ vary. We will also assume that the index $I=|\bf w|-d$ equals 1, as done in [16]. Since our weighted varieties produce Berglund-Hübsch duals embedded in non well-formed weighted projective spaces, in principal our procedure will give only upper bounds for the dimension of the moduli, however in most cases we found that these bounds are zero.
@@
The complex dimension of the moduli for $Z_{f^{T}}$ is bounded from below by
$\dim_{\mathbb{C}}H^{0}(\mathbb{P}(\tilde{\bf w}),\mathcal{O}(\tilde{d}))-\sum_{i}\dim_{\mathbb{C}}H^{0}(\mathbb{P}(\tilde{\bf w}),\mathcal{O}(\tilde{w}_{i}))$ (3.3.6)
We begin computing $)\dim_{\mathbb{C}}H^{0}(\mathbb{P}(\tilde{\bf w}),\mathcal{O}(\tilde{d}$)). Since this refers to the number of all possible monomials $z_{0}^{x_{0}}z_{1}^{x_{1}}z_{2}^{x_{2}}z_{3}^{x_{3}}z_{4}^{x_{4}}$ of degree $\tilde{d}$, we calculate this number counting the solutions of the Diophantine equation
@@
where the unknowns $x_{i}$ are non-negative integers such that at least one of them is non-zero. We will solve this equation adding certain mild constraints on the exponent $a_{1}$ and $m_{3}$: either $\gcd(a_{1},m_{3})=1$ and $a_{1}>2$ or $a_{1}=2.$ As we will see at the end of this section, the remaining cases of interest can be computed case by case.
@@
Let $f^{T}$ be an invertible polynomial described as in (3.3.3) with an associated weight vector $\tilde{\bf w}$ given as in (3.3.5). If we add the additional conditions $\gcd(a_{1},m_{3})=1$ and $a_{1}>2$, then the equation (3.3.7) has exactly five solutions. Moreover, the set of generators of $)H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(\tilde{d}$)) is given by the set
@@
    Since $\gcd(a_{1},m_{3})=1$, then we have $m_{3}\mid(m_{3}-\tilde{v}_{2}x_{2}-\tilde{v}_{3}x_{3}-\tilde{v}_{4}x_{4})$. That is, the expression $\tilde{v}_{2}x_{2}+\tilde{v}_{3}x_{3}+\tilde{v}_{4}x_{4}$ can assume two values: $m_{3}$ or 0. First, we suppose that $\tilde{v}_{2}x_{2}+\tilde{v}_{3}x_{3}+\tilde{v}_{4}x_{4}=m_{3}$. This forces $x_{1}=0$. Moreover, since $\tilde{v}_{2}$, $\tilde{v}_{3}$ and $\tilde{v}_{4}$ are given in (3.3.4) and these describe the weight vector $\tilde{\bf w}$ in (3.3.5), then the equation $\tilde{v}_{2}x_{2}+\tilde{v}_{3}x_{3}+\tilde{v}_{4}x_{4}=m_{3}$ has exactly three solutions: $(a_{2},1,0),(0,a_{3},1)$ and $(1,0,a_{4})$. Thus, we obtain three solutions for the equation (3.3.8): $(0,0,a_{2},1,0),(0,0,0,a_{3},1)$ and $(0,0,1,0,a_{4})$. Finally, if we suppose that $\tilde{v}_{2}x_{2}+\tilde{v}_{3}x_{3}+\tilde{v}_{4}x_{4}=0$, then $x_{1}=a_{1}$. In this case the only solution is $(0,a_{1},0,0,0)$.
@@
[…]
@@
This result follows from Lemma 3.3.1 and the fact that $\dim_{\mathbb{C}}H^{0}(\mathbb{P}(\tilde{\bf w}),\mathcal{O}(\tilde{w}_{i}))\geqslant 1$ for each $i$ and $\sum_{i=0}^{4}\dim_{\mathbb{C}}H^{0}(\mathbb{P}(\tilde{\bf w}),\mathcal{O}(\tilde{w}_{i}))\leqslant 5.$
@@
Let $f^{T}$ be an invertible polynomial described in (3.3.3) with associated weight vector $\tilde{\bf w}$ as in (3.3.5). If we add the additional condition $a_{1}=2$, then Equation (3.3.7) has six solutions. Moreover, the set of generators of $)H^{0}(\mathbb{P}(\textbf{w}),\mathcal{O}(\tilde{d}$)) is given by the set
@@
As $\gcd(m_{2},m_{3})=1$, then $m_{2}\mid(x_{0}+m_{2}x_{1})$. This implies that $m_{2}\mid x_{0}$. Moreover, from (3.3.9) we have $m_{3}x_{0}\leqslant 2m_{3}m_{2}$. Since $m_{3}\neq 0$, then $x_{0}\leqslant 2m_{2}$. Thus, $x_{0}$ can take three values: 0, $m_{2}$ or $2m_{2}$.
@@
        Nevertheless, the solutions of this last equation are $(a_{2},1,0),(0,a_{3},1)$ and $(1,0,a_{4})$, which implies that $2\hat{x}_{i}=1$ for any $i$. That is, $\hat{x}_{i}$ is not integer, which is not possible.
@@
        which has trivial solution. Therefore, the solution of Equation (3.3.9) is $(0,2,0,0,0)$.
@@
[…]
@@
First, we will prove that $\dim_{\mathbb{C}}H^{0}(\mathbb{P}(\tilde{\bf w}),\mathcal{O}(\tilde{w}_{1}))\geqslant 2$. In fact, it is equivalent to show that the quantity of solutions of the following Diophantine equation is not less that two:
@@
Here, we can exhibit at least two solutions: $(m_{2},0,0,0,0)$ and $(0,1,0,0,0)$. Therefore, we obtain $\dim_{\mathbb{C}}H^{0}(\mathbb{P}(\tilde{\bf w}),\mathcal{O}(\tilde{w}_{1}))\geqslant 2$.
On the other hand, we know that $\dim_{\mathbb{C}}H^{0}(\mathbb{P}(\tilde{\bf w}),\mathcal{O}(\tilde{w}_{i}))\geqslant 1$ for $i\neq 1$. This implies that
$6\leqslant\sum_{i=0}^{4}\dim_{\mathbb{C}}H^{0}(\mathbb{P}(\tilde{\bf w}),\mathcal{O}(\tilde{w}_{i}))\leqslant\dim_{\mathbb{C}}H^{0}(\mathbb{P}(\tilde{\bf w}),\mathcal{O}(\tilde{d}))=6.$
@@
$\dim_{\mathbb{C}}H^{0}(\mathbb{P}(\tilde{\bf w}),\mathcal{O}(\tilde{d}))-\sum_{i=0}^{4}\dim_{\mathbb{C}}H^{0}(\mathbb{P}(\tilde{\bf w}),\mathcal{O}(\tilde{w}_{i}))=0.$
@@
Proposition 3.3.3.
[…]
@@
    and the moduli of the orbifold $Z_{f^{T}}$ is included in the quotient
[…]
@@
  • If $f$ is a polynomial as in Lemma 3.3.2, then $)G(\mathbf{\tilde{w}}$)) is given on generators by
@@
    and the moduli of the orbifold $Z_{f^{T}}$ is included in the quotient
[…]
@@
$\square$
In [16], Theorem 4.1 we established the existence of 75 new rational homology 7-spheres admitting Sasaki-Einstein metrics with not well-formed quotients, that is, with some group elements having codimension 1 fixed point sets. All these links, which are listed in a table in the Appendix in [16], are given by polynomials of chain-cycle type of type II with weights satisfying the conditions given in (3.2.1). So we have:
Proposition 3.3.4.
Let $f$ be and invertible polynomial as in (3.3.1) with weights satisfying conditions (3.2.1). If its Berglund-Hübsch transpose $f^{T}$ with associated weight vector $\bf\tilde{w}$ satisfies the conditions given in either Lemma 3.3.1 or Lemma 3.3.2 and additionally $f$ belongs to one of the 236 rational homology 7-spheres admitting Sasaki-Einstein metrics found in [5] and [16], then the dimension of the local moduli space of Sasaki-Einstein metrics of the rational homology 7-spheres at $L_{f^{T}}$ equals zero. Thus rational homology 7-spheres given as links coming from polynomials $f^{T}$ as above do not admit inequivalent families of Sasaki-Einstein structures.
@@
From the list of 75 new examples of Sasaki-Einstein rational homology 7-sphere given in the Appendix of [16], there are 7 weight vectors $\bf\tilde{w}$ that do not satisfy the additional conditions of either Lemma 3.3.1 or Lemma 3.3.2, actually all of the elements of this table satisfy $a_{1}>2$ but $\gcd(m_{3},a_{1})\not=1.$ For these, we have the following table where $b_{\mu_{\mathbb{R}}}$ denotes the upper bound for the real dimension of the local moduli of Sasaki-Einstein metrics of rational homology 7-spheres at $L_{f^{T}}.$
$\bf\tilde{w}$ $\tilde{d}$ $m_{3}$ $)\dim_{\mathbb{C}}H^{0}(\mathbb{P}(\tilde{\bf w}),\mathcal{O}(\tilde{d}$)) $\sum_{i=0}^{4}\dim_{\mathbb{C}}H^{0}(\mathbb{P}(\tilde{\bf w}),\mathcal{O}(\tilde{w}_{i}))$ $b_{\mu_{\mathbb{R}}}$